OpenStudy (anonymous):

Find a - 2b + 3c diminished by a - 2b - 3c + d.

OpenStudy (anonymous):

i am going to guess that this means \[a-2b+3c-(a-2b-3c+d)\]

OpenStudy (anonymous):

remove parentheses using distributive law, combine like terms \[a-2b+3c-(a-2b-3c+d)=a-2b+3c-a+2b+3c-d\] \[=6c-d\]

OpenStudy (anonymous):

Thank you!